OpenStudy (maheshmeghwal9):

If a matrix is of order 2x2 & we have to prove its characteristic polynomial of order '3' ,then why does cayley-hamilton theorem does not apply on this question?

OpenStudy (maheshmeghwal9):

please help!

OpenStudy (maheshmeghwal9):

I mean \[|A-\lambda I|\] method.

OpenStudy (anonymous):

A n*n has order n polynomial so 2*2 matrix cannot have order 3 characteristic

OpenStudy (maheshmeghwal9):

but I have an ex.

OpenStudy (anonymous):

ex?!1

OpenStudy (maheshmeghwal9):

\[[A]=\left[\begin{matrix}2 & 3 \\ 1 & 2\end{matrix}\right]\] & this result holds true:- \[A^3-4A^2+A=0\]

OpenStudy (anonymous):

it is correct and clay hamilton work.try it again you multiply A to side of the equations.

OpenStudy (maheshmeghwal9):

K! thanx a lot !

OpenStudy (anonymous):

ok .are you study engineering ?

OpenStudy (maheshmeghwal9):

ya I am doing IIT-JEE coaching

OpenStudy (anonymous):

ok be successful

OpenStudy (maheshmeghwal9):

Thanx a lot! but what do u do?

OpenStudy (anonymous):

I also studied Control system Engineering.

OpenStudy (maheshmeghwal9):

oh I see ! good luck to u also.

OpenStudy (anonymous):

keep in touch i can help you so much!

OpenStudy (maheshmeghwal9):

k! thanx.

OpenStudy (anonymous):

your welcome